OpenStudy (anonymous):

How much money would have to be deposited per month (where the deposits will be made at the end of each month) at 6% interest compounded monthly to accumulate to $42,000 in 16 years.

OpenStudy (anonymous):

Your'e equation is : Interest = Principal × Rate × Time where: 'Interest' is the total amount of interest paid, 'Principal' is the amount lent or borrowed, 'Rate' is the percentage of the principal charged as interest each year. The rate is expressed as a decimal fraction, so percentages must be divided by 100. For example, if the rate is 15%, then use 15/100 or 0.15 in the formula. 'Time' is the time in years of the loan. www.teacherschoice.com

OpenStudy (anonymous):

Is this right? A=P(1+(r/n))^nt 42,000 = P(a+ (.06/12))^196 P = $16119.75 $16119.75/12 = 1343.31 Is $1343.31 correct?

OpenStudy (anonymous):

I meant to write 1 not "a" in the question.

OpenStudy (anonymous):

42000 =C(1+(.06/12)^ (12x16)) if you have a calculator - TI-83 you put 42000 into your y1 & the other side of the equation into your y2. & then find the intersecting points. yes you're right except for the number above its supposed to be 16x12 = 192 not 196

OpenStudy (anonymous):

Oh yes sorry about that typo. Wait how come when I check $1343.31 it's not giving me $42,000?

OpenStudy (anonymous):

I multiplied that with 16... I'm doing it wrong probably...

OpenStudy (anonymous):

your answer being : 16120.02 your calculator is going to say: x=16120.015 y=42000

OpenStudy (anonymous):

do you have a TI-83 graphing calculator?

OpenStudy (anonymous):

but isn't 16119.75 only for "Deposited now" and not "deposited per month"?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

did you get that answer? > 16119.75

OpenStudy (anonymous):

If you want a FINAL amount you keep your initial amount. If you do not you subtract it.

OpenStudy (anonymous):

Yea i didn't do it with the calculator though...I solved it using the formula because my teacher wants us to show work. So the final answer would be 1-the pvalue?

OpenStudy (anonymous):

If you're looking for your final amount, yes.